OpenStudy (anonymous):

Solve √x+3-3=x plz explain..

OpenStudy (anonymous):

Is the question \[\sqrt{x + 3} - 3 = x\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First, you bring -3 to the other side.\[\sqrt{x + 3} = x + 3\]Then, you square both sides.\[(\sqrt{x + 3})^{2} = (x + 3)^{2}\]Remember that this is binomial squared!\[x + 3 = x^{2} + 6x + 9\]Bring everything to one side.\[0 = x^{2} + 5x + 6\]Factor the quadratic.\[0 = (x + 3)(x + 2)\]Solve for x.\[x = -3, -2\]Now, plug this back into the original equation. Only one of them should work. The one that works is the correct answer.

OpenStudy (anonymous):

ty for your gud explaination. I have another question tho.

OpenStudy (anonymous):

when it comes out negative at the begining then it's no solution?

OpenStudy (anonymous):

What do you mean? Show the question so I know what you mean please :)

OpenStudy (anonymous):

-5√x-9-7=23.

OpenStudy (anonymous):

Yes. Here's an explanation. When you have a squared term and you are finding the square root to solve for the variable, you get a positive and negative answer. Example: \(x^{2} = 9 \rightarrow x = ±3\) However, if the square root is given to you and you are not solving with it, then you only get the positive answer. So in this case, you are correct and it's no solution.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np :)